# Exploring the Effect of Explanation Content and Format on User Comprehension and Trust

### v1

@@ 3.2. Explanations for the QCancer Algorithm
This definition thus allowed us to introduce two different formats of explanation. Firstly, chart-based occlusion-1 (OC) explanations are the explanations computed as in Definition 3.3 in chart format, as shown in Figure 2 with an accompanying textual description. We can see here that the features are the same as those selected by the SC explanations but the values are significantly different. However, the counterfactual reading of these explanations also allowed us to describe the attributions intuitively to users. Thus, we used the same occlusion-1 explanations, this time in textual format, to define text-based occlusion (OT) explanations, which are delivered as shown in Figure 3 with their accompanying textual description. This distinction between the two types of occlusion-1 explanation allows us to assess the effects not only of the content of an explanation, but also of its format. Note that SHAP allowed no such intuitive reading: the complexity of SHAP explanations in Definition 3.2 demonstrates they how do not lend themselves to text in what we might have called “ST” explanations, and that most users would not understand these game-theoretical notions. Indeed, it is difficult to picture, at least in a cognitively manageable form, what the ST equivalent for the SC explanation in Figure 1 would be, as the OT explanation in Figure 3 is to the OC explanation in Figure 2. We could have given an approximation of these explanations, but this simplification could be seen as being unrepresentative of SHAP, and it would be an unfair test. Thus, we refrained from defining and experimenting with this form of explanation.
@@
Figure 2. An OC explanation (bottom) for the vignette given in Example 3.1, and textual description of the OC explanation (top). Red (green) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
@@
Figure 3. An OT explanation (bottom) for the vignette given in Example 3.1, and textual description of the OT explanation (top). Note that the textual descriptions of the OT explanations were identical to those of the OC explanations.
3.3. Experimental Design & Procedure
@@
We assessed the RQs in two different samples: a sample from the general population where a limited number of participants had medical experience (Study 1) and an expert sample where participants had some medical training and experience (Study 2). We created two surveys using Qualtrics [^3] https://www.qualtrics.com that followed the same structure, as follows.
@@
  • Main Test (4 x 6 questions) – In the main task, we provided the participants with four vignette sets, each containing a vignette, a prediction and an explanation (with its accompanying textual description), along with six questions on the explanation. The four sets concerned two types of explanation for comparison, each with a low risk and a high risk vignette. The six questions consisted of:
    • an attention check comprising the most important feature task (attention check; free text variable)
    • a perceived comprehensibility rating (addressing RQ2; interval scale from 0 to 10)
    • a trust rating (addressing RQ3; interval scale from 0 to 10)
    • a request for reasoning for the trust rating (optional; free text variable);
    • a request for additional feedback (optional; free text variable)
    • a comprehensibility test comprising the definition recognition task (addressing RQ1; categorical variable with four options).
@@
After each vignette, the participants were provided with an the attention check to assess whether they were engaging with the experiment. The attention check consisted in asking them which was the most important feature in determining the gastro-oesophageal cancer risk according to the explanation, with the correct answer being the feature which gave the highest absolute value of $e_{S}$ for SC explanations and $e_{O}$ for OC and OT explanations. For the perceived comprehensibility rating addressing RQ2 and the trust rating addressing RQ3, we asked the participants how well they thought they understood and how much they would trust, respectively, the explanation, on a scale of 0-10. Next, we asked for free-form textual feedback from the participants on reasons for their trust rating and on any additional points, both of which were optional. Finally, on a new page, we measured the participants’ comprehension of the explanation, addressing RQ1, by offering them four differently worded descriptions of the explanation as indicated below, one of which was correct for each explanation (as indicated in square brackets, but not shown to participants), and asked them which they believed it to be correct.
@@
$\bullet$ The change in predicted risk when the feature’s current value is modified to the baseline value. [Correct for OC and OT explanations.]
$\bullet$ A representative importance of this feature’s value relative to the baseline value, based on different configurations of values for other features. [Correct for SC explanations.]
$\bullet$ The portion of the predicted risk associated with that feature.
@@
Figure 4 illustrates the experimental design of Studies 1 and 2, which were identical in this regard. The participants were split such that 50% assessed SC vs. OC explanations (i.e. addressing RQ1a, RQ2a and RQ3a) and 50% assessed SC vs. OT explanations (i.e. addressing RQ1b, RQ2b and RQ3b). We then split the twenty vignettes into one vignette for demonstration to all participants, ten vignettes presenting low risk patients (less then 2% chance of gastro-oesophageal cancer) and nine vignettes presenting high risk patients (equal to or greater than 2% chance of gastro-oesophageal cancer), to ensure that all participants received one low and one high risk for each of the explanations that they examined. We then further split the high and low risk sets into two (giving High Risk Set A, High Risk Set B, Low Risk Set A and Low Risk Set B) so that the vignettes were presented to the participants evenly. Further, the four vignette sets for each participant were presented in random order to ensure there was no ordering bias.
Figure 4. Experimental design for user studies, with the two sets of high risk vignettes represented by shades of red, and the two sets of low risk vignettes represented by shades of green.
3.4. Participant Recruitment and Ethics Statement
For Study 1, we aimed to recruit 100 participants from the general population, who were over 18 and spoke English, via the online research platform Prolific [^4] Prolific (https://www.prolific.com) was chosen because of its superior data quality when compared with other platforms (Peer et al., 2022).. Upon completion, the participants were paid an average of \$4.50 for their time based on the platform’s standard reimbursement rates of \$21.20 per hour. Data collection was anonymous during Study 1. In Study 2, we aimed to recruit at least 50 students in medicine from UK-based universities. The potential participants were approached via student mailing lists, and were invited via a link to the survey in an email, which also contained the participant information sheet with details about the study. Note that all students were 18 years of age and English-speaking since they were attending university in the UK. At the end of Study 2, participants were asked to provide their university email address, to which we sent a £10 voucher for the online retailer Amazon to reimburse them for their time. All the participant responses were pseudononymised once this payment was made (i.e. we deleted their e-mail addresses from the dataset). Participation in both studies was voluntary, participants were asked to provide informed consent after reading the participant information sheet and they could withdraw at any time prior to completing the survey, simply by closing the questionnaire window, with their data deleted permanently. We informed the participants that once they had completed the survey, they could no longer withdraw and we would keep their data. [^5] The anonymised and pseudononymised data (respectively) will be shared via Zenodo (https://zenodo.org) on acceptance. We received ethics approval from our institution for these studies.
3.5. Data Analysis
Creation of the Main Variables of Interest
@@ 3.5. Data Analysis
Statistical Analyses
We ran the analyses in R (version 4.3.1). We built generalised linear and linear mixed-effects regression models with the lme4 R package (Bates et al., 2015) to analyse our categorical and interval variables, respectively. For the generalised linear regression models, we could only include random intercepts by participants (the models did not converge with more random intercepts), and, for the linear regression models, we included random intercepts by participants and vignettes. For the generalised linear and the linear regression models we reported Odds Ratios (ORs) and raw regression slopes (bs), respectively. We also reported 95% CIs for each regression coefficient. To test our hypotheses, we used the conventional p-value threshold of .05.
4. Results
@@
Study 1 was completed by 99 participants [^6] One participant skipped the survey in Study 1 so we deleted their data and did not include them in the results., from which 1 reported to have colour-blindness (we did not measure demographics). Study 2 was completed by 69 medical students [^7] One participant left a non-university email in Study 2 so we deleted their data and did not include them in the results., from which 3 reported to have colour-blindness. The mean age of the participants in Study 2 was 22.3 (SD = 3.3), with 35 females (32 males, 1 other and 1 preferred not to disclose). As expected, participants in the medical student sample (Study 2) reported to have heard more about medical algorithms than participants in the Prolific sample (Study 1), but the number of people who had tried a medical AI tool before was negligible in both samples. Interestingly, trust towards algorithms in general was comparable in the two samples, and, on average, participants in both samples had neutral attitudes (roughly 5 on a scale from 1 to 9, where 5 was neutral). Table 2 presents the descriptive statistics broken down by samples. Since our sample sizes were defined by resource constraints, we conducted sensitivity power analyses as per the recommended good practice (Lakens, 2022). The multilevel regressions directly comparing the SC with either the OC or the OT explanations (RQ2 and RQ3) were powered at the 80% level to detect effect sizes of Cohen‘s d of 0.31 and 0.38, in Study 1 and 2, respectively. [^8] We used GPower 3.1 for the sensitivity power analyses, and we calculated the effective sample sizes for each study by multiplying the number of participants by the cluster sizes (i.e. each user responded to two vignettes) and then dividing it by the Design effect, which is 1+ (cluster size - 1) multiplied by the intra-class correlation (which we assumed to be low, r=0.2). Hence, we applied a Design effect of 1.2 for our studies.
Variable Response Study 1 (N = 99) Study 2 (N = 69)
Familiarity with Unfamiliar 85 (86%) 41 (59%)
AI (count) Not used 12 (12%) 26 (38%)
Used 2 (2%) 2 (3%)
Mean 4.84 5.20
Trust (0-10) Median 5 5
SD 2.17 2.00
Table 2. Descriptive statistics of the familiarity and trust in medical AI variables broken down by study samples.
@@
We analysed the data from the most important feature task to assess whether the participants engaged with the experiment. Since all of our low risk vignettes used the same scaling for the chart-based explanations, our SC explanations were visualised with systematically smaller bars than the OC explanations, which may have created unintended difficulty in differentiating the bars on the charts for the former. Therefore, we calculated the percentages of correctly identified most important features separately for low and high risk cases, as well as for the explanations (See Table 3). Generally, the participants managed to correctly pinpoint the most important features in both studies, except for the chart-based explanations of the low risk cases in Study 1, where they failed in around half of the instances. These findings are in line with the notion of participants understanding and engaging with the task at hand but struggling to correctly respond when the task is more difficult.
Explanation High Risk Low Risk
Study 1 Study 2 Study 1 Study 2
SC 80.8% 92.8% 45.5% 69.6%
OC 85.1% 84.4% 57.4% 87.5%
OT 88.5% 94.6% 90.3% 89.2%
Table 3. Percentages of correctly identified most important features (Attention Check).
@@
We merged the OC and OT explanations to a single occlusion explanation group, since format was not relevant for the definition recognition task. First, we ran two intercept only regression models to compare performance in recognising the SHAP and occlusion definitions against chance level. In all cases, except for the occlusion explanations in Study 2, the 95% CIs of the ORs included chance level (OR of 1/3) suggesting that the participants struggled to recognise the correct definition of the explanations. Next, we analysed all data in a single regression model and entered explanation type as a predictor. We found no evidence that participants were better at recognising the definition of occlusion than SHAP explanations. Table 4 presents the corresponding ORs, their 95% CIs and the multilevel regressions.
Regression Model Study 1 Study 2
OR p OR p
SHAP 0.48 [0.28, 0.80] - 0.32 [0.12, 0.87] -
Occlusion 0.33 [0.16, 0.68] - 0.64 [0.38, 1.08] -
SHAP vs. Occlusion 0.88 [0.57, 1.36] .578 1.27 [0.76, 2.13] .359
Table 4. Multilevel regression models predicting the objective comprehensibility of explanations (the definition recognition task for RQ1).
@@
In both studies, the reported level of subjective understanding was higher for the OT (Study 1: Mean (M) = 6.57, Standard Deviation (SD) = 2.37; Study 2: M = 6.73, SD = 2.21) than for the SC explanations (Study 1: M = 5.67, SD = 2.51; Study 2: M = 6.18, SD = 2.62). In comparison, the level of reported understanding of the OC explanations was comparable to that of the SC explanations in Study 1 (OC Study 1: M = 5.62, SD = 2.27), and higher than the reported understanding of SC explanations in Study 2 (OC Study 2: M = 6.81, SD = 2.13). The analyses revealed that when SC explanations were compared to all occlusion explanations, participants perceived the latter to be significantly more comprehensible than the former in both studies. However, the comparison of the different formats of the occlusion explanations to the SC explanations in separate analyses only found evidence for the superiority of OT over SC explanations. Moreover, in Study 1, participants found OT more comprehensible than OC explanations. The other comparisons were non-significant. Table 5 presents the results of the multilevel regression analyses.
Regression Model Study 1 Study 2
b p b p
 SC vs. OC & OT  0.46 [0.18, 0.74]  .001 0.58 [0.18 - 0.98]  .004
SC vs. OC 0.10 [-0.30, 0.49]  .637 0.51 [-0.07, 1.09]  .088
SC vs. OT 0.79 [0.41, 1.17]  ¡ .001  0.67 [0.18, 1.16]  .008
OC vs. OT 0.95 [0.10, 1.79]  .030 -0.09 [-0.97, 0.79]  .841
Table 5. Multilevel regression models predicting the difference in subjective understanding of explanations (RQ2).
@@
The reported level of trust was consistently the lowest for the SC explanations across the studies (Study 1: M = 4.95, SD = 2.45; Study 2: M = 6.73, SD = 2.21). The OT explanations scored the highest in Study 1 (Study 1: M = 5.47, SD = 2.32; Study 2: M = 5.66, SD = 2.18), and the OC explanations scored the highest in Study 2 (Study 1: M = 5.20, SD = 2.27; Study 2: M = 6.03, SD = 1.97). Similarly to the analyses of the subjective understanding of the explanations, when the occlusion explanations were compared with the SC explanations, participants trusted the occlusion explanations significantly more in both studies. However, the follow-up analyses revealed some contradictions between the two studies. Namely, Study 1 showed evidence for the superiority of OT over SC explanations, while Study 2 showed evidence for the preference for OC over SC explanations in terms of their trustworthiness. The other comparisons remained non-significant. Table 6 presents the results of the multilevel regression analyses.
Regression Model Study 1 Study 2
b p b p
SC vs. OC & OT 0.40 [0.13, 0.66]  .003 0.38 [0.02, 0.75]  .042
SC vs. OC 0.33 [-0.07, 0.73]  .107 0.57 [0.04, 1.10]  .035
SC vs. OT 0.44 [0.10, 0.78]  .012 0.23 [-0.25, 0.71]  .345
OC vs. OT 0.27 [-0.56, .1.10]  .530 0.37 [-0.48, 1.21]  .397
@@
Table 6. Multilevel regression models predicting the difference in the trustworthiness of explanations (RQ3).
5. Discussion and Limitations

### v2

@@ 3 Methodology
This definition thus allowed us to introduce two different formats of explanation. Firstly, chart-based Occlusion-1 (OC) explanations are the explanations computed as in Definition 2 in chart format, as shown in Figure 2 with an accompanying textual description. We can see here that the features are the same as those selected by the SC explanations but the values are significantly different. Further, the counterfactual reading of these explanations also allowed us to describe the attributions intuitively to users. Thus, we used the same Occlusion-1 explanations, this time in textual format, to define text-based Occlusion-1 (OT) explanations, which are delivered as shown in Figure 3 with their accompanying textual description. This distinction between the two types of Occlusion-1 explanation allows us to assess the effects not only of the content of an explanation, but also of its format. [^4] SHAP allowed no such intuitive reading: the complexity of SHAP in Definition 1 demonstrates they how do not lend themselves to text in what we might have called “ST” explanations, and that most users would not understand these game-theoretical notions. Indeed, it is difficult to picture, at least in a cognitively manageable form, what the ST equivalent for the SC explanation in Figure 1 would be, as the OT explanation in Figure 3 is to the OC explanation in Figure 2. We could have given an approximation of these explanations, but this simplification could be seen as being unrepresentative of SHAP, and it would be an unfair test. Thus, we refrained from defining and experimenting with this form of explanation.
@@
Figure 2: OC explanation (bottom) for the vignette given in Example 1, and textual description thereof (top). Red (green) bars represent features which increased (decreased, respectively) the risk of cancer. The patient’s values are given in bold and the baseline values in grey.
@@
Figure 3: OT explanation (bottom) for the vignette given in Example 1, and textual description thereof (top). Note that the textual descriptions are identical for the OC and OT explanations.
Experimental Design and Procedure
@@
We assessed the RQs in two different samples: one from the general population where a limited number of participants had medical experience, representing stakeholders in the form of patients (Study 1), and an expert sample where participants had some medical training and experience, representing stakeholders in the form of healthcare practitioners (Study 2). We created two Qualtrics [^5] https://www.qualtrics.com surveys using the following structure (see the Supplementary Material for an example):
@@
  • Main Test (4 x 6 questions) – In the main task, we provided the participants with four vignette sets, each containing a vignette, a prediction and an explanation (with its textual description), along with six questions on the explanation. The four sets concerned two types of explanation for comparison, each with a low risk and a high risk vignette. The six questions consisted of:
    • an attention check comprising the most important feature task (attention check, free text variable);
    • a perceived comprehensibility rating (addressing RQ2, interval scale from 0 to 10);
    • a trust rating (addressing RQ3, interval scale from 0 to 10);
    • a request for reasoning for the trust rating (optional, free text variable);
    • a request for additional feedback (optional; free text variable);
    • a comprehensibility test comprising the definition recognition task (addressing RQ1, categorical variable with four options).
@@
After each vignette, the participants were provided with an the attention check to assess whether they were engaging with the experiment. The attention check consisted in asking participants which was the most important feature in determining the gastro-oesophageal cancer risk according to the explanation, with the correct answer being the feature which gave the highest absolute value of $e_{S}$ for SC explanations and $e_{O}$ for OC and OT explanations. For the perceived comprehensibility rating (RQ2) and the trust rating (RQ3), we asked the participants how well they thought they understood and how much they would trust, respectively, the explanation, on a scale of 0-10. Next, we asked for free-form textual feedback from the participants on reasons for their trust rating and on any additional points, both of which were optional. Finally, on a new page, we measured the participants’ comprehension of the explanation (RQ1) by offering four differently worded descriptions of the explanation as indicated below, one of which was correct for each explanation (as indicated in square brackets, but not shown to participants), and asked them which one they believed to be correct.
@@
$\bullet$ The change in predicted risk when the feature’s current value is modified to the baseline value. [Correct for OC/OT.]
$\bullet$ A representative importance of this feature’s value relative to the baseline value, based on different configurations of values for other features. [Correct for SC.]
$\bullet$ The portion of the predicted risk associated with the feature.
@@
Figure 4 illustrates the experimental design of Studies 1 and 2, which were identical in this regard. The participants were split such that 50% assessed SC versus OC explanations (i.e. addressing RQ1a, RQ2a and RQ3a) and 50% assessed SC versus OT explanations (i.e. addressing RQ1b, RQ2b and RQ3b). We then split the twenty vignettes into one vignette for demonstration to all participants, ten vignettes presenting low risk patients and nine vignettes presenting high risk patients (as defined earlier), to ensure that all participants received one low and one high risk for each of the explanations that they examined. We then further split the high and low risk sets into two (giving High Risk Sets A and B, and Low Risk Sets A and B) so that all vignettes were presented evenly. Further, the four vignette sets for each participant were presented in random order to ensure there was no ordering bias.
Figure 4: Experimental design for both user studies, with the two sets of high risk vignettes represented by shades of red, and the two sets of low risk vignettes represented by shades of green.
Participant Recruitment
For Study 1, we aimed to recruit 100 participants from the general population, who were over 18 and spoke English via the online research platform Prolific [^6] Prolific (https://www.prolific.com) was chosen for its superior data quality when compared with other platforms Peer et al. (2022).. Upon completion, the participants were paid an average of \$4.50 for their time based on the platform’s standard reimbursement rates of \$21.20 per hour. Data collection was anonymous during Study 1. In Study 2, we aimed to recruit at least 50 students in medicine from UK-based universities. The potential participants were approached via student mailing lists, and were invited via a link to the survey in an email, which also contained the participant information sheet with details about the study. All students were at least 18 years of age and English-speaking as they were attending university in the UK. At the end of Study 2, participants were asked to provide their university email address, to which we sent a £10 voucher for the online retailer Amazon to reimburse them. [^7] The anonymised and pseudononymised data will be shared via Zenodo (https://zenodo.org) on acceptance.
Data Analysis
@@
We ran the statistical analyses in R (version 4.3.1). We built generalised linear and linear mixed-effects regression models with the lme4 R package Bates et al. (2015) to analyse our categorical and interval variables, respectively. For the generalised linear regression models, we only included random intercepts by participants, and for the linear regression models, we included random intercepts by participants and vignettes. For the generalised linear and the linear regression models we report Odds Ratios (ORs) and raw regression slopes (bs), respectively. We also report 95% CIs for each regression coefficient. To test our hypotheses, we used a p-value of 0.05.
We ran two types of analyses for RQ2 and RQ3. First, we compared the SHAP to a single Occlusion-1 explanation group, merging OC and OT explanations, for a general test of explanation content. Second, we made pairwise comparisons of explanations to examine the effects of explanation content, format and their interaction. We did not conduct formal analyses of the free-text responses of general comments and the justifications about trust towards the explanations due to the low response rates (15% across questions and both studies).
4 Results
@@
Study 1 was completed by 99 participants, from which 1 reported to have colour-blindness (we did not measure demographics). Since 5 participants failed the attention check, we did not include their data in any of our analyses (see more details below) resulting in a final sample of 94. Study 2 was completed by 69 medical students, from which 3 reported to have colour-blindness. [^8] One participant skipped the survey in Study 1 and one participant left a non-university email in Study 2; we deleted the data of both participants and did not include them in the results. The mean age of the participants in Study 2 was 22.3 (SD = 3.3), with 35 females (32 males, 1 other and 1 preferred not to disclose). As expected, participants in the medical student sample (Study 2) reported to have heard more about medical algorithms than participants in the Prolific sample (Study 1), but the number of people who had tried a medical AI tool before was negligible in both samples (Study 1: Unfamiliar = 85% [80/94], Not used = 13% [12/94], Used = 2% [2/94]; Study 2: Unfamiliar = 59% [41/69], Not used = 38% [26/69], Used = 3% [2/69]). Interestingly, trust towards algorithms in general was comparable in the two samples, and, on average, participants in both samples had neutral attitudes (measured on a scale of 0 to 10; Study 1: Mean (M) = 4.78, Standard Deviation (SD) = 2.18; Study 2: M = 5.20, SD = 2.00). Since our sample sizes were defined by resource constraints, we conducted sensitivity power analyses as per the recommended good practice Lakens (2022). The multilevel regressions directly comparing the SC with the OC or OT explanations (RQ2 and RQ3) were powered at the 80% level to detect effect sizes of Cohen‘s d of 0.31 and 0.38, in Study 1 and 2, respectively. [^9] We used GPower 3.1 for the sensitivity power analyses, and we calculated the effective sample sizes for each study by multiplying the number of participants by the cluster sizes (i.e. each user responded to two vignettes) and then dividing it by the design effect, which is 1+ (cluster size - 1) multiplied by the intra-class correlation (which we assumed to be low, r=0.2). Hence, we applied a design effect of 1.2 for our studies.
@@
We analysed the data from the most important feature task to assess whether the participants engaged with the experiment. Generally, the participants were able to correctly pinpoint the most important features in both studies. In Study 1, 72%, whereas in Study 2, 86% of the participants identified at least 3 out of the 4 most important features correctly. However, in Study 1, 5 participants failed to identify any of the 4 most important features, therefore, we did not include their data in any of the analyses.
@@
We merged the OC and OT explanations to a single Occlusion-1 explanation group, since format was not relevant for the definition recognition task. First, we ran two intercept only regression models to compare performance in recognising the SHAP and Occlusion-1 definitions against chance level. The OR of recognising the correct SHAP definition remained around the chance level (OR of 1/3) in both studies (Study 1: OR = 0.46 [0.24, 0.76], Study 2: OR = 0.32 [0.12, 0.87]). In contrast, the OR of recognising the correct Occlusion-1 definition was substantially above the chance level in Study 2 (Study 1: OR = 0.31 [0.10, 0.60], Study 2: OR = 0.64 [0.38, 1.08]). We then analysed all data in a single regression model and used explanation type as a predictor. We found no evidence that participants were better at recognising Occlusion-1 over SHAP explanations (Study 1: OR = 1.14 [0.73, 1.78], p = .566; Study 2: OR = 1.27 [0.76, 2.13], p = .359).
@@
In both studies, the reported level of subjective understanding was higher for the OT (Study 1: (M) = 6.55, (SD) = 2.39; Study 2: M = 6.73, SD = 2.21) than for the SC explanations (Study 1: M = 5.62, SD = 2.51; Study 2: M = 6.18, SD = 2.62). In comparison, the level of reported understanding of the OC explanations was comparable to that of the SC explanations in Study 1 (OC Study 1: M = 5.49, SD = 2.28), and higher than the reported understanding of SC explanations in Study 2 (OC Study 2: M = 6.81, SD = 2.13). The analyses revealed that when SC explanations were compared to all Occlusion-1 explanations, participants perceived the latter to be significantly more comprehensible than the former in both studies. However, the comparison of the different formats of the Occlusion-1 explanations to the SC explanations in separate analyses only found evidence for the superiority of OT over SC explanations. Moreover, in Study 1, participants found OT more comprehensible than OC explanations. The other comparisons were non-significant. Table 2 presents the results of the multilevel regression analyses.
Regression Study 1 Study 2
Model b p b p
 S v O  0.45 [0.16, 0.73]  .001  0.58 [0.18, 0.98]  .004
SC v OC 0.04 [-0.37, 0.45]  .847  0.51 [-0.07, 1.09]  .088
SC v OT  0.80 [0.42, 1.19]  .001  0.67 [0.18, 1.16]  .008
OC v OT   1.06 [0.19, 1.94]   .030   -0.09 [-0.97, 0.79]   .841
Table 2: Multilevel regression models for the difference in subjective understanding of SHAP and Occlusion-1 explanations (RQ2).
@@
The reported level of trust was consistently the lowest for the SC explanations across the studies (Study 1: M = 4.82, SD = 2.39; Study 2: M = 6.73, SD = 2.21). The OT explanations scored the highest in Study 1 (Study 1: M = 5.41, SD = 2.33; Study 2: M = 5.66, SD = 2.18), and the OC explanations scored the highest in Study 2 (Study 1: M = 5.06, SD = 2.24; Study 2: M = 6.03, SD = 1.97). Similarly to the analyses of the subjective understanding of the explanations, when the Occlusion-1 explanations were compared with the SC explanations, participants trusted the Occlusion-1 explanations significantly more in both studies. However, the follow-up analyses revealed some contradictions between the two studies. Namely, Study 1 showed evidence for the superiority of OT over SC explanations, while Study 2 showed evidence for the preference for OC over SC explanations in terms of their trustworthiness. The other comparisons remained non-significant. Table 3 presents the results of the multilevel regression analyses.
Regression Study 1 Study 2
Model b p b p
S v O 0.44 [0.17, 0.70]  .003  0.38 [0.02, 0.75]   .042
SC v OC 0.39 [-0.03, 0.80]  .066  0.57 [0.04, 1.10]   .035
 SC v OT  0.46 [0.12, 0.81]   .009  0.23 [-0.25, 0.71]   .345
OC v OT 0.36 [-0.50, 1.22]  .414  0.37 [-0.48, 1.21]   .397
@@
Table 3: Multilevel regression models for the difference in the trustworthiness of SHAP and Occlusion-1 explanations (RQ3).
5 Discussion